\begin{thebibliography}{99}

\bibitem{gse}
Gashkov S. B., Sergeev I. S. On the complexity of linear Boolean
operators with thin matrices. J. Applied and Industrial Math.
2011. {\bf5}(2), 202--211.

\bibitem{gs1e}
Grinchuk M. I., Sergeev I. S. Thin circulant matrices and lower
bounds on the complexity of some Boolean operators. Diskretn.
Anal. Issled. Oper. 2011. {\bf 18}(5), 38--53 (in Russian).

\bibitem{jue}
Jukna S. Boolean function complexity. Advances and frontiers.
Springer--Verlag, 2012.

\bibitem{lue}
Lupanov O. B. On rectifier and switching-and-rectifier circuits.
Dokl. Akad. Nauk SSSR. 1956. {\bf 111}(6), 1171--1174 (in
Russian).

\bibitem{she}
Sch\"{o}nhage A. Schnelle multiplikation von polynomen \"{u}ber
k\"{o}rpern der charakteristik 2. Acta Inf. 1977. {\bf 7},
395--398.

\bibitem{toe}
Toom A. L. The complexity of a scheme of functional elements
realizing the multiplication of integers. Soviet Math. Doklady.
1963. {\bf 4}, 714--716.

\end{thebibliography}

\documentclass[12pt,a4paper]{article}
\usepackage{latexsym,amssymb,amsmath,amsthm}

\newtheorem{theorem}{Theorem}

\begin{document}
\title{Implementation of linear maps with circulant matrices via modulo 2 rectifier circuits of bounded depth\footnote{Research supported in part by RFBR, grants
11--01--00508, 11--01--00792, and OMN RAS ``Algebraic and
combinatorial methods of mathematical cybernetics and information
systems of new generation'' program (project ``Problems of optimal
synthesis of control systems'').}}
\date{}
\author{Igor S. Sergeev\footnote{e-mail: isserg@gmail.com}}

\maketitle

\begin{abstract}
In the present note we show that for any constant $k \in \mathbb
N$ an arbitrary Boolean circulant matrix can be implemented via
modulo~2 rectifier circuit of depth $2k-1$ and complexity
$O\left(n^{1+1/k}\right)$, and also via circuit of depth $2k$ and
complexity $O\left(n^{1+1/k}\log^{-1/k} n\right)$.
\end{abstract}

Recall that {\it rectifier $(m,n)$-circuit} is an oriented graph
with $n$ vertices labeled as inputs and $m$ vertices labeled as
outputs. {\it Modulo $2$ rectifier circuit} implements a Boolean
$m\times n$ matrix $A=(a_{i,j})$ iff for any $i$ and $j$ the
number of oriented paths from $j$-th input to $i$-th output is
congruent to $a_{i,j}$ modulo~2. Complexity of a circuit is the
number of edges in it, circuit depth is the maximal length of an
oriented path. See details in~\cite{jue,lue}.

$n\times n$ matrix $Z=(z_{i,j})$ is {\it circulant} iff for any
$i$, $j$ one has $z_{i,j} = z_{0,k}$, where $k=(j-i)\bmod n$.

Consider a linear map with Boolean circulant $n\times n$
matrix~--- it computes a cyclic (algebraic, over $GF(2)$)
convolution with some constant vector $A$. Indeed, components of
vector $C=(C_0,\ldots,C_{n-1})$ which is a convolution of vectors
$A=(A_0,\ldots,A_{n-1})$ and $B=(B_0,\ldots,B_{n-1})$ satisfy
formulae:
$$ C_k = \sum_{i+j \equiv k \mod n} A_iB_j. $$

The following theorem allows to extend results~\cite{gse,gs1e} on
comparison of complexity of implementation of some circulant
matrices via rectifier circuits and modulo 2 rectifier circuits to
bounded depth circuits.

\begin{theorem}
For any $k \in \mathbb N$ an arbitrary Boolean circulant $n\times
n$ matrix $Z$ can be implemented via modulo $2$ rectifier circuit:

a) of depth $2k-1$ and complexity at most $f(2k-1)n^{1+1/k}$;

b) of depth $2k$ and complexity at most $f(2k)n\left(\frac{n}{\log
n}\right)^{1/k}$.
\end{theorem}

The proof is by induction. For $k=1$ we use a trivial depth-1
circuit of complexity $O(n^2)$ and a circuit of depth 2 and
complexity $O(n^2/\log n)$ provided by O.B. Lupanov's
method~\cite{lue}.

Now we prove an induction step from $k-1$ to $k$. We use the
polynomial multiplication method due to A.L. Toom~\cite{toe}
together with À.~Sch\"onhage's idea~\cite{she} allowing to extend
the method to binary polynomials. The depth-$d$ polynomial
multiplication is reduced to several parallel depth-$(d-2)$
multiplications.

Split a vector of variables into $q$ blocks of length $n/q$ and
interpret each block as a vector of coefficients of a polynomial
from the ring
$$ R=GF(2)[y]/(y^{2\cdot3^s}+y^{3^s}+1).$$
Parameter $s$ satisfies condition $3^s \ge n/q$.

So, multiplication of binary polynomials of degree $n-1$ can be
performed as a multiplication of polynomials of degree $n/q-1$
over $R$. The latter multiplication can be performed via DFT of
order $3^m \ge 2q$ with primitive root $\zeta=y^{s+1-m}\in R$.

Next, we describe a circuit.

Its input is a polynomial $B(x)=\sum B_i x^i \in R[x]$ of degree
$q-1$. A constant factor is denoted by $A(x)=\sum A_i x^i$. Output
is the product $C(x)=A(x)B(x)=\sum C_i x^i$.

1. Compute $B(\zeta^0)$, \ldots, $B(\zeta^{3^m-1})$.

2. Compute $C(\zeta^i)=A(\zeta^i)B(\zeta^i)$ for all
$i=0,\ldots,3^m-1$.

3. Compute coefficients of $C(x)$.

We implement stages 1 and 3 via depth-1 circuits and stage 2 ---
via circuit of depth $d-2$. Next, we estimate the circuit
complexity, denote it by $M(d,n)$.

1. Multiplication by a power of $y$ in $R$ has linear complexity.
Hence, the value of polynomial $F(x)$ at the point $y^p$ can be
computed with linear complexity as well. Therefore, the complexity
of stage 1 is $O(3^m3^sq)$.

2. Every multiplication at the stage 2 is a multiplication of
binary polynomials of degree $2\cdot3^s-1$ with a subsequent
modulo reduction. Perform multiplication via circuit of depth
$d-2$ and complexity $M(d-2,2\cdot3^s)$ provided by induction
hypothesis. Reduction of a polynomial $g(y)$ (here, of degree
$4\cdot3^s-1$) modulo $y^{2\cdot3^s}+y^{3^s}+1$ is performed via
duplication of some of its coefficients (that is, outputs of a
preceding subcircuit) and identifying of some coefficients, since
every coefficient of $g$ is to be used at most twice. Henceforth,
modulo reduction can be embedded into multiplication circuit with
no depth increasing and with at most doubling of the circuit
complexity. Thus, the total complexity of stage 2 is at most
$2\cdot3^mM(d-2,2\cdot3^s)$.

3. By the fundamental property of DFT, coefficients of $C(x)$
satisfy $C_i=C^*(\zeta^{-i})$, where polynomial $C^*(x)$ has
coefficients $C(\zeta^i)$. Thus, the complexity of stage 3 is that
of stage 1, $O(3^m3^sq)$.

To transform the product of polynomials over $R$ backward to the
product of binary polynomials, one performs a substitution
$x=y^{2\cdot3^s}$. The substitution preserves depth and complexity
of the circuit.

The choice of parameters to obtain required complexity bounds is:
$q=n^{1/k}$ for $d=2k-1$ and $q=(n/\log n)^{1/k}$ for $d=2k$;
$3^s=\Theta(n/q)$, $3^m=\Theta(q)$.

(By construction, $f(k)=O(c^k)$ for some constant $c$.) \qed

\end{document}